OpenStudy (anonymous):

HOW DO I SOLVE THIS MEDAL Donna teaches dance in groups of five students. She rents a hall for the dance classes at $450 per month. She charges $40 per hour per student. Donna teaches four groups of students, and her net earnings per month are $4,350. Part A: If each student learns dance for the same number of hours, write an equation that can be used to calculate the number of hours each student learns dance in a month. (5 points) Part B: How many hours did each student learn dance each month? Show your work and justify each step. (5 points)

OpenStudy (anonymous):

i think its 450x+40(20)=4350

OpenStudy (anonymous):

for part a

OpenStudy (anonymous):

hmm well I'm not sure of the answer just yet. But in your equation I don't think the 450x is right. First of all she pays 450 a month so you would want to subtract that from however much money she receives. Secondly, The x is not necessary, We're only dealing with one month here so whats the point in the x?

OpenStudy (anonymous):

okay so what should i do

OpenStudy (anonymous):

Ok, I'm thinking it should be in this form \[20(40x)-450=4350\]

OpenStudy (anonymous):

Does that make sense to you?

OpenStudy (anonymous):

i know that theres 4 group of kids and 5 in each so thats 20 kids and she charges 40 for each kid so that would be 800 dollars in total for all the kids

OpenStudy (anonymous):

yeah it makes sense

OpenStudy (anonymous):

Right. It would be 40 per kid if they all only danced one hour. We right the x to represent number of hours.

OpenStudy (anonymous):

Write*

OpenStudy (anonymous):

ok im a little confused so the answer for hours would be 6

OpenStudy (anonymous):

Thats what I'm getting. Sounds like a reasonable number to me.

OpenStudy (anonymous):

okay thanks do you think you could me with some other questions

OpenStudy (anonymous):

I can certainly give them a shot.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

The table below shows the outputs y for different inputs x: Input (x) 1 1 3 3 Output (y) 4 2 12 6 Part A: Does the table represent y as a function of x? Justify your answer. (5 points) Part B: The total cost f(x), in dollars, for renting a paddleboat for x hours is shown below: f(x) = 20 + 10x What is the value of f(120), and what does f(120) represent? (5 points)

OpenStudy (anonymous):

Ok, so in order for something to be considered a function it must pass the vertical line test. Mean that if we put a vertical line on the function at any point it should only go through the function one time. More simply put, if its a function we should have one one y value for every x value. Think about a circle. A circle is NOT a function. If we were to put a vertical line on the function we would get two y values for every x value.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So based on that, would the given points be part of a function?

OpenStudy (anonymous):

so then it doesnt represent y as a function

OpenStudy (anonymous):

Right. Because at x=1 we have two y values. Same goes for x=3.

OpenStudy (anonymous):

got it so y does not represent x as a function because it doesnt pass through a vertical line

OpenStudy (anonymous):

Pretty much. If you're writing an answers a more correct way of putting it would be that the relation does not pass the vertical line test.

OpenStudy (anonymous):

A relation is what we call a graph that isn't a function by the way.

OpenStudy (anonymous):

ok and for part b what do i do

OpenStudy (anonymous):

Part b is super easy. When you're dealing with a function in the form F(x)=9x+23 if you want to find say F(4) just plug 4 in where ever you see an x in the function.

OpenStudy (anonymous):

okay and for part c What is the value of f(120), and what does f(120) represent? what do i do there

OpenStudy (anonymous):

okay so im guessing i would do f(120)=20 + 10x and just solve it

OpenStudy (anonymous):

so the answer would be 10 hrs

OpenStudy (anonymous):

Right. Plugging 120 in for every x in the function. As far as what it represents, the question pretty much tells you that. It represents how much is would cost to rent a boat for 120 hours.

OpenStudy (anonymous):

ok thanks now this is the second question

OpenStudy (anonymous):

Well for one your answer would be an amount of money. It might be helpful to reread the question.

OpenStudy (anonymous):

Here, I'll show you how its worked out.

OpenStudy (anonymous):

actually im a little confused i put f(120)=20+10x I know that f(120) is the total cost so when i solve it i get 10 hrs for x

OpenStudy (anonymous):

is x supposed to equal f(x) like the same number

OpenStudy (anonymous):

\[f(x)=20+10x----> f(120)=20+(10)(120)----->f(120)=20+1200---->f(120)=1220\]

OpenStudy (anonymous):

Oh it got cut off.

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(120)=20+1200---->f(120)=1220\]

OpenStudy (anonymous):

so x = 10

OpenStudy (anonymous):

So you aren't solving for x. All you need to do is plug in 120 for every x in the function. Remember x is equal to the number of hours the boat was rented. Well we were told in the question that that was 120 right?

OpenStudy (anonymous):

cuz the question says that f(120) is the total cost and for 20+10x 20 is the fee and 10 is how much they charge per hour so wouldnt it be 10 hrs

OpenStudy (anonymous):

cuz 10 times 10 is 100 plus the 20 would equal 120 right

OpenStudy (anonymous):